Edward is 5 times as old as Timothy TODAY. In 5 years Edward will be 4 times as old as Timothy. How old is Edward?
Answer by stanbon(75887)

Edward is 5 times as old as Timothy TODAY. In 5 years Edward will be 4 times as old as Timothy. How old is Edward?
----
Equations:
E = 5T
----
E+5 = 4(T+5)
------
Substitute for "E" and solve for "T":
5T + 5 = 4T + 20
------
T = 15 (Tim's age now)
------
Solve for "E":
E = 5T
E = 5*15 = 75 yrs (Edward's age now)
--------------------------------------
Cheers,
Stan H.
====================